### v1

@@ Abstract
This paper presents a state-of-the-art algorithm for the vertex enumeration problem of arrangements, which is based on the proposed new pivot rule, called the Zero rule. The Zero rule possesses several desirable properties: i) It gets rid of the objective function; ii) Its terminal satisfies uniqueness; iii) We establish the if-and-only-if condition between the Zero rule and its valid reverse, which is not enjoyed by earlier rules; iv) Applying the Zero rule recursively definitely terminates in $d$ steps, where $d$ is the dimension of input variables. Because of so, given an arbitrary arrangement with $v$ vertices of $n$ hyperplanes in $\mathbb{R}^{d}$, the algorithm’s complexity is at most $\mathcal{O}(n^{2}d^{2}v)$ and can be as low as $\mathcal{O}(nd^{4}v)$ if it is a simple arrangement, while Moss’ algorithm takes $\mathcal{O}(nd^{2}v^{2})$, and Avis and Fukuda’s algorithm goes into a loop or skips vertices because the if-and-only-if condition between the rule they chose and its valid reverse is not fulfilled. Systematic and comprehensive experiments confirm that the Zero rule not only does not fail but also is the most efficient.
@@
Hyperplane arrangement; vertex enumeration; polynomial pivot rule.
@@ 1 Introduction
Vertex enumeration (VE) is a fundamental problem identifying vertices that satisfy specific linear constraints [5, 9, 15, 16, 26, 35]. When these constraints are linear inequalities like $Ax\leq b$, it entails enumerating vertices on a polytope, while constraints such as $Ax=b$ correspond to enumerating vertices on an arrangement. Geometrically, VE bridges two crucial representations of a polytope or an arrangement: $H$-representation and $V$-representation, which respectively define a polytope or an arrangement from its faces and vertices. As such, VE is not only a fundamental problem in computational discrete geometry but also holds a wide array of real-world applications. VE has been applied in problems including, among others, effective field theory [39], implicit mesh generation [23], and computing approximating polytopes [3]. More recently, VE has emerged as a key component in unraveling the inner workings of deep learning [17, 37, 38]. The underlying concept stems from the fact that a network with piecewise linear activation functions serves as a piecewise linear entity, partitioning the space into numerous linear regions. Each region is defined by a set of linear constraints. Identifying the vertices of these linear constraints is pivotal in understanding the dynamics and robustness of deep networks [38].
@@
The earliest published method to solve the problem of VE can be found in [26]. Their approach is to add inequalities in a stepwise manner. In each step, “new” extreme points are created and “old” ones are excluded. Subsequently, due to the intrinsic connection between VE and linear programming, the simplex method and its variants, such as Abdullahi and Chen’s algorithm [1], Avis and Fukuda’s pivoting method [4], Balinski’s algorithm [6], Bremner’s primal simplex method [9], Chand and Misra’s algorithm [10], Chvatal’s linear programming approach [11], Dyer and Frieze’s complexity analysis [15], Elbassioni and Raychaudhuri’s enumeration algorithm [16], Mattheiss’s approach [24], Provan’s efficient algorithm [28], Shioura’s optimal algorithm [31], were utilized to solve the VE problem. The simplex method was originally used in linear programming, based on the observation that a linear objective function always attains its maximum at extreme points of the feasible region (polytope). Geometrically, one only needs to move the current vertex to an adjacent one, and if the maximum exists, the algorithm will terminate at one vertex that maximizes the objective function. Specifically, in this method, a dictionary is used to represent the current vertex, and a step called “pivoting” is employed to select entering and leaving variables to update the dictionary, thus traversing between neighboring vertices. However, if a vertex lies on more than $d$ hyperplanes, the pivoting may trap into several dictionaries concerning the same vertex, which greatly deteriorates the efficiency. Hence, different pivot rules were introduced to avoid this situation and boost efficiency. For example, the Bland’s rule [7, 8], the Edmonds-Fukuda rule [18] and its variant [12], the Criss-Cross rule [32, 33, 36] and its variant [19], and the Jensen’s general relaxed recursion [21].
@@
Many previous papers were dedicated to the VE of a polytope. Although the importance of VE of an arrangement is increasingly recognized, to the best of our knowledge, only a few methods focus on VE of an arrangement. The method in Moss [25] is designed for listing exactly one co-basis from each orbit under the action from a set of automorphisms to the arrangement. It considers the smallest positive value and the largest negative value in the ratio test of every column, thereby obtaining all adjacent dictionaries. By repeating this process, every dictionary is obtained. Then, the algorithm compares each dictionary in this arrangement to determine whether it is a duplicate. This method is essentially a brute-force enumeration, with a high demand for storage and computational complexity. The method proposed by Avis and Fukuda [4] is based on an observation that since repeating one pivot rule can yield a path from some dictionary to an optimal dictionary, then tracking back the path can generate all dictionaries ending at that optimal dictionary, without the need for the brute-force enumeration. Instead of directly computing coordinates of all vertices, the algorithm outputs lexicographically minimum basis for each vertex, and then coordinates can be computed based on the outputs if needed.
@@
Despite the attraction of the idea, the algorithm in Avis and Fukuda [4] should be upgraded from the following three major aspects: i) The chosen Criss-Cross rule and its valid reverse, unfortunately, do not satisfy a necessary and sufficient (if-and-only-if) condition, which fails the searching because it makes the pivoting either go to a loop or catastrophically skip some vertices. ii) The objective function is not necessary for the VE of an arrangement. Firstly, the objective function restricts the traversal of vertices, leading to more steps in complex arrangements. Secondly, the value of the objective function is unpredictable until reaching a vertex, requiring considering all vertices to optimize the objective function, which is more time-consuming. Thirdly, the pivot operation on a dictionary can be performed without the objective function, and neither the magnitude of the objective function nor the objective function itself affects our computation of the vertex coordinates based on the dictionary’s basis or co-basis. Therefore, the objective function only serves as a guide during the pivot process, but in our study, this guidance can also be observed from the dictionary itself. iii) The Criss-Cross rule may not have a unique convergence, which introduces more operations in the VE problem and makes the algorithm difficult to organize.
In this paper, building upon the framework proposed in Avis and Fukuda [4], we develop a new pivot rule that is specifically tailored for the VE of an arrangement. This rule is not a small modification but solves the key issues of the Criss-Cross rule and leads to a more efficient algorithm. The new pivot rule, dubbed as the Zero rule, uses the linear independence of normal vectors among hyperplanes in the arrangement. The most distinct features of the Zero rule in contrast to the Criss-Cross rule are that i) the if-and-only-if condition between the Zero rule and its valid reverse is established, which can ensure all vertices are discovered; ii) it gets rid of objective functions; iii) its convergence is unique, and the number of convergence steps is bounded by the dimension $d$, which arranges all vertices in a single and shallow tree, thereby saving computation. Given a simple arrangement with $v$ vertices of $n$ hyperplanes in $\mathbb{R}^{d}$, the complexity of the algorithm using the Zero rule is at most $\mathcal{O}(n^{2}d^{2}v)$ but can be as low as $\mathcal{O}(nd^{4}v)$ for simple arrangements. Table 1 summarizes the characteristics of Moss [25], Avis and Fukuda [4], and our proposed algorithm. Furthermore, we provide systematic and comprehensive experiments to compare the performance of different algorithms. Experiments confirm that our proposed algorithm using the Zero rule not only does not fail but also is the most efficient.
Table 1: A summary between Moss [25], Avis and Fukuda [4] and ours, where $n$ is number of hyperplanes, $d$ is the dimension, and $v$ is the number of vertices.
storage obj. func. no cycle complexity single tree
[25] $\mathcal{O}(ndv)$ $\mathcal{O}(nd^{2}v^{2})$
[4] $\mathcal{O}(nd)$
Ours $\mathcal{O}(1)$ $\mathcal{O}(nd^{4}v)$ or $\mathcal{O}(n^{2}d^{2}v)$
@@ 2 Notations and Definitions
Definition 2.1 (Hyperplane and arrangement).
[…]
Definition 2.2 (Vertex of an arrangement).
A vertex of an arrangement is the unique solution to the system of $d$ equations corresponding to $d$ intersecting hyperplanes. Moreover, a vertex is degenerate if it is contained in more than $d$ hyperplanes.
The vertex enumeration problem for a hyperplane arrangement is enumerating all vertices of the arrangement.
Definition 2.3 (Dictionary with the objective function).
Suppose there are $n$ hyperplanes in an arrangement, denoted by slack variables $x_{i}=b_{i}x_{g}-\langle\mathbf{c}_{i},\mathbf{y}\rangle,i=1,\cdots,n$, where $x_{g}$ has a constant value $1$. Let $x_{f}$ be the objective function which is the linear combination of elements in $\mathbf{y}$. For a set of linearly independent vectors $\{\mathbf{c}_{j_{1}},\cdots,\mathbf{c}_{j_{d}}\}$, $y_{1},\cdots,y_{d}$ can be expressed in terms of $\mathbf{x}_{j_{1}},\cdots,\mathbf{x}_{j_{d}}$, leading to a relation between slack variables:
@@
where $\mathbf{x}_{B}=(x_{i_{1}},\cdots,x_{i_{n-d}},x_{f})$ and $\mathbf{x}_{N}=(x_{g},x_{j_{1}},\cdots,x_{j_{d}})$. Such a relation is called a dictionary. A dictionary is degenerate if it is associated with a degenerate vertex. Figure 1 shows a dictionary with the objective function.
Definition 2.4 (Basis and co-basis).
In a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, the set $B=\{i_{1},\cdots,i_{n-d},f\}$ is called basis, and $\mathbf{x}_{B}$ are basic variables. $N=\{g,j_{1},j_{2},\cdots,j_{d}\}$ is called co-basis, and $\mathbf{x}_{N}$ are non-basic variables. $\bar{A}=(\bar{a}_{ij})\in\mathbb{R}^{{(n-d+1)}\times{(d+1)}}$ is the coefficient matrix, and $\bar{a}_{ij}$ expresses entries in $\bar{A}$ corresponding to $x_{i}$ and $x_{j}$.
Definition 2.5 (Dictionary without the objective function).
As Figure 1 shows, a dictionary can go without the objective function. Accordingly, simply removing the last row of Eq. (1), $\mathbf{x}_{B}=(x_{i_{1}},\cdots,x_{i_{n-d}})$ and $B=\{i_{1},\cdots,i_{n-d}\}$.
In the proposed Zero rule, all dictionaries work without the objective function. Each dictionary is associated with one vertex, while each non-degenerate vertex is associated with only one dictionary. The vertex enumeration problem can be transformed into enumerating all dictionaries. The key in this process is to circumvent yielding vertices associated with multiple dictionaries more than once.
Definition 2.6 (Primal & Dual Feasible [4]).
In a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, if $\bar{a}_{ig}\geq 0$, then the variable $x_{i},i\in B_{\neq f}$ is primal feasible. If $\bar{a}_{fj}\leq 0$, then the variable $x_{j},j\in N_{\neq g}$ is dual feasible. A dictionary is primal feasible, if $\forall i\in B_{\neq f},x_{i}$ is primal feasible. A dictionary is dual feasible, if $\forall j\in N_{\neq g},x_{j}$ is dual feasible. The dictionary is optimal, if it is both primal and dual feasible. Note that these definitions are only applicable to the dictionary with the objective function.
@@
Definition 2.7 (Pivoting).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary, a pivot $(r,s)$ on the dictionary is an exchange between $x_{r}$ in $\mathbf{x}_{B}$ and $x_{s}$ in $\mathbf{x}_{N}$, which will generate a new dictionary $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$, where the new coefficient matrix $\tilde{A}=(\tilde{a}_{ij})$ is obtained by computing Eq. (2) and sorting rows and columns of $\tilde{A}=(\tilde{a}_{ij})$ in increasing order of indices:
@@
Definition 2.8 (Valid reverse pivoting).
Let $D$ be an arbitrary dictionary, a reverse pivot $(s,r)$ on $D$ will return a new dictionary $\tilde{D}$. It is a valid reverse about the rule $\mathcal{R}$ if applying $\mathcal{R}$ on $\tilde{D}$ fixed at $(s,r)$ instead of other entries.
Definition 2.9 (Lexicographically minimum basis [4]).
Let $B_{1},B_{2},\cdots,B_{n}$ be basis for all dictionaries associated with the same vertex. The basis attaining the minimum lexicographic order among $B_{1},B_{2},\cdots,B_{n}$ is called the lexicographically minimum basis. Verifying the lexicographically minimum basis and only computing the vertex using this dictionary can avoid producing the same vertex more than once.
It is important to note that for a non-degenerate vertex associated with only one dictionary, this dictionary (non-degenerate dictionary) is always lexicographically minimum. Testing if the basis of a degenerate dictionary is lexicographically minimum is called the lex-min test. Let $B$ be a basis for a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$. $B$ is not lexicographically minimum if and only if there exists $r\in B_{\neq f}$ and $r\in N_{\neq g}$ such that $r>s,\bar{a}_{rg}=0$ and $\bar{a}_{rs}\neq 0$ [4].
The lex-min test can avoid outputting a vertex more than once: First, given a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, let $B_{1}=\{i\in B_{\neq f}~{}|~{}\bar{a}_{ig}=0\}$, which selects the indices whose associated constant columns are zero. Then, $x_{i}=0$ for $i\in B_{1}$, which means that all the $i$-th hyperplanes, $i\in B_{1}\cup N$, pass through the same vertex. Consequently, the $i$-th hyperplanes, $i\notin B_{1},i\in B$ will never pass through this vertex. Thus, checking rows covered by $B_{1}$ is sufficient to analyze, without the need for really solving the dictionary. If a dictionary $D$ is not lex-min (e.g. LHS of Figure 2), then there will always exist another dictionary $\tilde{D}$ in the entire arrangement (e.g. RHS of Figure 2), where the lexicographical order of $\tilde{B}$ is smaller than $B$. Since the lexicographically smallest dictionary is unique at each vertex, we can ensure that each vertex is counted only once, thereby saving the computational overhead.
Figure 2: An example of the lex-min test. LHS: not pass the test. RHS: pass the test.
Definition 2.10 (The Criss-Cross rule).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary with an objective function. The Criss-Cross rule [32, 33, 36] is as follows: 1) $i\neq f,g$ is the smallest index such that $x_{i}$ is primal or dual infeasible. 2) If $i\in B_{\neq f}$, let $r=i$ and $s$ be the minimum index such that $\bar{a}_{rs}>0$; otherwise, let $s=i$ and let $r$ be the minimum index such that $\bar{a}_{rs}<0$. 3) Pivot $(r,s)$ and update the dictionary.
The condition to verify valid reversibility of the Criss-Cross rule that was proposed in [4] is that if $(s,r)$, where $s\in B_{\neq f}$ and $r\in N_{\neq g}$, is a valid reverse of the Criss-Cross rule for a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, then either (a) $\bar{a}_{sg}>0,\bar{a}_{sr}>0,\bar{a}_{sj}$ for $j\in N_{\neq g},j<s$, or (b) $\bar{a}_{fr}<0,\bar{a}_{sr}<0,\bar{a}_{ir}<0~{}\mathrm{for}~{}i\in B_{\neq f},i<r$.
Theorem 2.11.
The condition provided by Avis and Fukuda in [4] is a necessary but not sufficient condition of valid reversibility of the Criss-Cross rule.
Proof: Assuming that from the dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, by pivoting $(s,r)$, we obtain $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$. The condition (a) ensures the selection of $s$ in the second step of the Criss-Cross rule, which gives a specific row $r\in\tilde{B}$. But it cannot determine whether the first step of the Criss-Cross rule on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ identifies the desired $r$-th row. Similarly, the condition (b) guarantees the selection of $r$ in the second step of the Criss-Cross rule, which is given by a specific row $s\in\tilde{N}$. However, it cannot guarantee whether the first step of the Criss-Cross rule will yield the required column $r$. Therefore, even if $(s,r)$ in $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ satisfies their given conditions, it still cannot guarantee that applying the Criss-Cross rule on $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ will pinpoint $(r,s)$.
$\Box$

### v2

@@ Abstract
This paper presents a state-of-the-art algorithm for the vertex enumeration problem of arrangements. We introduce a new pivot rule, called the Zero rule. The Zero rule possesses several desirable properties: i) It eliminates the objective function; ii) Its terminal dictionary is single; iii) We establish the if-and-only-if condition between the Zero pivot and its valid reverse pivot; iv) Applying the Zero pivot recursively definitely terminates in $d$ steps, where $d$ is the dimension of the input variables. Based on this rule and its properties, we leverage it to obtain a more efficient vertex enumeration algorithm. Theoretically, given an arrangement in $\mathbb{R}^{d}$ composed of $n$ hyperplanes with $v$ vertices, where $v_{d}$ of those vertices reach the terminal in exactly $d$ steps when applying the Zero pivot, the algorithm’s complexity is $\mathcal{O}(n^{2}d^{2}(v-v_{d})-ndv_{d})$. For the class of simple arrangements, the complexity is $\mathcal{O}(nd^{4}v)$, which significantly improves $\mathcal{O}(n^{2}dv)$ of the Avis and Fukuda algorithm. Systematic and comprehensive experiments confirm that our algorithm is effective.
@@
[…]
@@ 1 Introduction
Vertex enumeration (VE) is a fundamental problem aimed at identifying vertices that satisfy specific linear constraints [3, 9, 13, 15, 30, 42]. When these constraints take the form of linear inequalities, such as in the condition $Ax\leq b$, it involves enumerating vertices on a polytope, while the condition $Ax=b$ corresponds to enumerating vertices on an arrangement. Geometrically, VE bridges two crucial representations of a polytope or an arrangement: $H$-representation and $V$-representation, which respectively define a polytope or an arrangement from its faces and vertices. As such, VE is not only a fundamental problem in computational discrete geometry but also holds a wide array of real-world applications. VE has been applied in problems including, among others, effective field theory [48] and computing approximating polytopes [1]. More recently, VE has emerged as a key component in unraveling the inner workings of deep learning [5, 16, 33, 34, 44, 47]. The underlying concept stems from the fact that a network with piecewise linear activation functions serves as a piecewise linear entity, partitioning the space into numerous linear regions. Each region is defined by a set of linear constraints. Identifying the vertices of these linear constraints is pivotal in understanding the dynamics and robustness of deep networks [47].
@@
The earliest published method to solve the problem of VE can be found in [30]. Their approach is to add inequalities in a stepwise manner. In each step, "new" extreme points are created and "old" ones are excluded. Subsequently, due to the intrinsic connection between VE and linear programming, the simplex method and its variants, such as Avis and Fukuda’s pivoting method [2], Balinski’s algorithm [4], Bremner’s primal simplex method [9], Chand and Misra’s algorithm [10], Chvatal’s linear programming approach [11], Dyer and Frieze’s complexity analysis [13] , Mattheiss’s approach [27], Dyer, Martin E, and Proll’s method [14] were utilized to solve the VE problems. The simplex method was originally used in linear programming, based on the observation that a linear objective function always attains its maximum at extreme points of the feasible region (polytope). Geometrically, one only needs to move the current vertex to an adjacent one, and if the maximum exists, the algorithm will terminate at one vertex that maximizes the objective function. Specifically, in this method, a dictionary is used to represent the current vertex, and a step called "pivoting" is employed to select entering and leaving variables to update the dictionary, thus traversing between neighboring vertices. However, if a vertex lies on more than $d$ hyperplanes, the pivoting may trap into several dictionaries concerning the same vertex, which greatly deteriorates the efficiency. Hence, different pivot rules were introduced to avoid this situation and boost efficiency. For example, the Bland’s rule [7, 8], the Criss-Cross rule [38, 39, 43] and its variant [17], the Jensen’s general relaxed recursion [22], the Random-Facet pivot rule and its variants [20, 23, 25], the Random-Edge pivot rule [19, 26], and the Facet pivot rule [45, 46].
@@
Many previous papers were dedicated to the VE of a polytope. Although the importance of VE of an arrangement is increasingly recognized, to the best of our knowledge, only a few methods focus on it: The method in Moss [29] is designed for listing exactly one co-basis from each orbit under the action from a set of automorphisms to the arrangement. It considers the smallest positive value and the largest negative value in the ratio test of every column, thereby obtaining all adjacent dictionaries, and hence all vertices. By repeating this process, every dictionary is obtained. Then, the algorithm compares each dictionary in this arrangement to determine whether it is a duplicate. This method is essentially a brute-force enumeration, with a high demand for storage and computational complexity. The method proposed by Avis and Fukuda [2] is based on an observation that since repeating one pivot rule can yield a path from some dictionary to an optimal dictionary, then tracking back the path can generate all dictionaries ending at that optimal dictionary, without the need for the brute-force enumeration. Instead of directly computing coordinates of all vertices, the algorithm first outputs lexicographically minimum basis for each vertex, and then computes coordinates.
@@
Despite the attraction of the idea, the algorithm in Avis and Fukuda [2] could be upgraded from the following three major aspects: i) The method for examining valid reverse pivots takes several steps, which is less efficient than an explicit if-and-only-if condition according to our analysis in Section 4.3 and Appendix. ii) The objective function is not necessary for the vertex enumeration of an arrangement. Firstly, the objective function restricts the traversal of vertices, resulting in more steps in complex arrangements. Secondly, the value of the objective function is unpredictable until a vertex is reached, which means that the objective function cannot be optimized until visiting all vertices. This is more time-consuming. Thirdly, the pivot operation on a dictionary can be performed without the objective function, and neither the magnitude of the objective function nor the function itself affects computing the vertex coordinates based on the dictionary’s basis or co-basis. Therefore, the objective function merely guides the pivot process, however, this guidance can be directly captured from the dictionary itself. iii) Multiple optimal vertices or dictionaries may exist in the Criss-Cross rule, which introduces additional steps and storage.
In this paper, we first introduce a pivot rule that can be applied throughout the hyperplane arrangement, referred to as the Zero rule, which utilizes the linear independence of the normal vectors of hyperplanes in the arrangement. Compared to the Criss-Cross rule, its most notable features are: i) the number of Zero pivots from each dictionary to the terminal is upper bounded by the dimension $d$, which is a small constant bound; ii) it eliminates the objective function; iii) the collection of dictionaries from which the Zero rule cannot select a suitable position is always unique, while the Criss-Cross rule may have multiple such dictionaries. In addition, we also give an if-and-only-if condition for detecting if a pivot is a valid reverse Zero pivot.
Based on these desirable properties of the Zero rule, we realize that the key points of the algorithm proposed by Avis and Fukuda [2] can be further improved with the help of the Zero rule. Compared to the algorithm of Avis and Fukuda [2], using our algorithm can effectively reduce both space and time complexity. Specifically, given a hyperplane arrangement in $\mathbb{R}^{d}$ consisting of $n$ distinct hyperplanes with $v$ vertices, there are $\mathcal{O}(v)$ dictionaries in this arrangement. Note that applying the Zero pivot takes at most $d$ iterations. Therefore, we denote the number of dictionaries that require exactly $d$ iterations as $\mathcal{O}(v_{d})$. Then the computational complexity of the algorithm improved by the Zero rule is $O(n^{2}d^{2}(v-v_{d})+ndv_{d})$. If a hyperplane arrangement is a simple arrangement [41], where $v=\binom{n}{d}$ and $v_{d}\geq\binom{n-d}{d}$, the complexity favorably turns into at most $\mathcal{O}(nd^{4}v)$, which significantly improves Moss [29] and Avis & Fukuda [2]. Moreover, since there is no additional storage, the needed space are for those variables generated during the computation, which are at most two dictionaries, in the order of $\mathcal{O}(nd)$. Table 1 summarizes the information of the proposed algorithm and its counterparts.
storage obj. func. complexity single tree
Ours $\mathcal{O}(nd)$ $\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})$
AF [2] $\mathcal{O}(nd)$ $\mathcal{O}(n^{2}dv)$
Enhanced AF $\mathcal{O}(nd)$ $\mathcal{O}(n^{2}dv)$
Moss [29] $\mathcal{O}(ndv)$ $\mathcal{O}(nd^{2}v^{2})$
Table 1: A summary between the proposed algorithm, Enhanced AF, Avis and Fukuda’s algorithm (AF [2]), Moss’ algorithm [29] (Moss). In our algorithm, $v_{d}$ is the number of dictionaries that require exactly $d$ steps of pivoting. If a hyperplane arrangement is a simple arrangement, the complexity of our algorithm is $\mathcal{O}(nd^{4}v)$, while for a hyperplane arrangement where $v_{d}$ accounts for a small portion, the complexity is $\mathcal{O}(n^{2}d^{2}v)$. Fortunately, the simple arrangement is quite general.
The structure of this paper is as follows: Section 2 introduces the definitions and notations used throughout this paper. Section 3 elaborates on the important properties of the proposed Zero rule. Section 4 includes: 1) a brief description of the algorithm proposed by Avis and Fukuda [2]; 2) the VE algorithm use the Zero Rule and advantage analysis, which is our algorithm; 3) a discussion of the complexities of various vertex enumerating algorithms. Section 5 presents detailed toy examples. Section 6 covers systematic experiments. Section 7 concludes this paper. In addition to the main content, we also give an enhanced version to the AF algorithm for independent interest to some readers in the Appendix, we denote the enhanced version as Enhanced AF.
@@ 2 Notations and Definitions
Definition 1 (Hyperplane and arrangement).
A hyperplane in $\mathbb{R}^{d}$ is defined as
$\{y\in\mathbb{R}^{d}:\langle\mathbf{c},\mathbf{y}\rangle=c_{1}y_{1}+c_{2}y_{2}+\ldots+c_{d}y_{d}=b,~{}\mathbf{c}\in\mathbb{R}^{d}\}$. The finite collection of hyperplanes is called hyperplane arrangement.
Definition 2 (Vertex of an arrangement).
A vertex of an arrangement is the unique solution to the system of $d$ equations corresponding to $d$ intersecting hyperplanes. Moreover, a vertex is degenerate if it is contained in more than $d$ hyperplanes. A vertex is also called a basic solution in linear programming.
In this paper, we assume that the arrangement contains at least one vertex. Particularly, simple arrangements are those hyperplane arrangements such that every vertex is non-degenerate and every $d$ hyperplanes form a vertex. This definition can also be found on page 530 of the book [41].
Definition 3 (Slack variable).
For each hyperplane, $\langle\mathbf{c}_{i},\mathbf{y}\rangle=b_{i}$, the corresponding slack variable is defined as $x_{i}=b_{i}x_{g}-\langle\mathbf{c}_{i},\mathbf{y}\rangle$, where $x_{g}$ is a constant-valued variable that is always equal to $1$.
It is important to note that if there are $d$ hyperplanes in the arrangement, $\langle\mathbf{c}_{j_{i}},y_{0}\rangle=b_{j_{i}},~{}i=1,\cdots,d$, such that their coefficients $\{\mathbf{c}_{j_{1}},\cdots,\mathbf{c}_{j_{d}}\}$ are linearly independent, then by solving the system of equations $x_{j_{i}}=b_{j_{i}}x_{g}-\langle\mathbf{c}_{j_{i}},\mathbf{y}\rangle$, we can express $\mathbf{y}$ in terms of $x_{g}$ and those slack variables. In this case, if we set $x_{j_{i}}=0,~{}\forall i=1,\ldots,d$, meaning that the corresponding $\mathbf{y}$ lies on all $d$ hyperplanes simultaneously, then the point $\mathbf{y}$ is a basic solution (vertex).
Definition 4 (Dictionary with an objective function).
Using the above notations, assume there are $n$ hyperplanes in an arrangement. Let $x_{f}$ be the objective function, which is a linear combination of elements in $\mathbf{y}$. If $\{\mathbf{c}_{j_{1}},\cdots,\mathbf{c}_{j_{d}}\}$ is linearly independent, express $y_{1},\cdots,y_{d}$ in terms of $x_{g},x_{j_{1}},\cdots,x_{j_{d}}$. Consequently, put $\{i_{1},\cdots,i_{n-d}\}=\{1,\cdots,n\}\backslash\{j_{1},\cdots,j_{d}\}$, by plugging this representation into the objective function and other slack variables, we have
@@
where $\mathbf{x}_{B}=\{x_{i_{1}},\cdots,x_{i_{n-d}},x_{f}\}$, $\mathbf{x}_{N}=\{x_{g},x_{j_{1}},\cdots,x_{j_{d}}\}$. Such an equation system is called a dictionary. Sometimes we also denote a dictionary by $D$ or $(B,N,\bar{A})$.
Definition 5 (Dictionary and Basic solution).
Using the above notations, setting $x_{j_{1}},\cdots,x_{j_{d}}$ to zeros induces a basic solution $\mathbf{y}_{0}$, we say $\mathbf{y}_{0}$ be corresponding basic solution of the dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$. In particular, if $\mathbf{y}_{0}$ is a degenerate vertex, we say $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ a degenerated dictionary.
Definition 6 (Basis and co-basis).
In a dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$, $\mathbf{x}_{B}$ are basic variables, the collection of their indices, $B=\{i_{1},\cdots,i_{n-d},f\}$ is called basis. $\mathbf{x}_{N}$ are non-basic variables, the collection of their indices, $N=\{g,j_{1},j_{2},\cdots,j_{d}\}$ is called co-basis. $\bar{A}=(\bar{a}_{ij})$ is the coefficient matrix, and $\bar{a}_{ij}$ expresses entries in $\bar{A}$ corresponding to $x_{i}\in\mathbf{x}_{B}$ and $x_{j}\in\mathbf{x}_{N}$.
In some case, we need to consider basic variables without the objective function or nonbasic variables without $x_{g}$. Thus, we put $B_{\neq f}=B\backslash\{f\}=\{i_{1},\cdots,i_{n-d}\}$ and $N_{\neq g}=N\backslash\{g\}=\{j_{1},j_{2},\cdots,j_{d}\}$.
Definition 7 (Dictionary without the objective function).
By removing the row of $x_{f}$ in Eq. (1), we say it is a dictionary without objective function. In this case, $\mathbf{x}_{B}=(x_{i_{1}},\cdots,x_{i_{n-d}})$ and $B=B_{\neq f}=\{i_{1},\cdots,i_{n-d}\}$.
@@
Definition 8 (Primal Feasible).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary. If $\bar{a}_{ig}\geq 0$, for some $i\in B_{\neq f}$, then the variable $x_{i}$ is primal feasible. A dictionary is primal feasible, if $x_{i}$ is primal feasible $\forall i\in B_{\neq f}$,.
Definition 9 (Dual Feasible and Optimal).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary with the objective function. If $\bar{a}_{fj}\leq 0$ for some $j\in N_{\neq g}$, then the variable $x_{j}$ is dual feasible. A dictionary is dual feasible, if $x_{j}$ is dual feasible $\forall j\in N_{\neq g}$. The dictionary is optimal, if it is both primal and dual feasible.
Definition 10 (Pivoting rule).
The pivoting rule is a method of selecting a pair of variables on a dictionary for subsequent operations.
Definition 11 (The Criss-Cross rule).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary with an objective function. The Criss-Cross rule [38, 39, 43] selects the variables $(r,s)$ in the following order:
1) $i\neq f,g$ is the smallest index such that $x_{i}$ is primal or dual infeasible.
2) If $i\in B_{\neq f}$, let $r=i$ and $s$ be the minimum index such that $\bar{a}_{rs}>0$; otherwise, let $s=i$ and let $r$ be the minimum index such that $\bar{a}_{rs}<0$.
If the step 1)1)1 ) does not apply, then the dictionary is optimal.
Definition 12 (Pivot).
Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary, a pivot $(r,s)$ on the dictionary is an exchange between $x_{r}$ in $\mathbf{x}_{B}$ and $x_{s}$ in $\mathbf{x}_{N}$, which generates a new dictionary $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$, where the new coefficient matrix $\tilde{A}=(\tilde{a}_{ij})$ is obtained by computing Eq. (2) and sorting rows and columns of $\tilde{A}=(\tilde{a}_{ij})$ in increasing the order of indices:
@@
In particular, if the pair $(r,s)$ on the dictionary $D$ is obtained by the $\mathcal{R}$ rule, we call the pivot $(r,s)$ an $\mathcal{R}$ pivot.
In fact, the pivot $(r,s)$ on a dictionary is moving $x_{s}$ to the left hand side and $x_{r}$ to the right hand side in $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$. This equation is to update the coefficient matrix and then rearrange the variables in a new basis and co-basis in lexicographic order to create a new dictionary.
Definition 13 (Valid Reverse Pivot).
Let $D$ be a dictionary obtained by a pivot $(r,s)$ on another dictionary $\tilde{D}$. Then the pivot $(s,r)$ on $D$ is called a valid reverse pivot. In particular, if the pivot $(r,s)$ on the dictionary $\tilde{D}$ is a $\mathcal{R}$ pivot, the pivot $(s,r)$ on $D$ is called a valid reverse $\mathcal{R}$ pivot.
According to the previous formulation, the vertex enumeration problem can be transformed into enumerating all dictionaries. The key in this process is to circumvent yielding vertices associated with multiple dictionaries more than once. Section 4 gives the method that employs the reverse idea for vertex enumeration.